A rectangle has a perimeter of 84cm.If 4cm is taken from the length and added to the breadth it becomes a square.Find the dimensions and area of the rectangle.
Answer by ramkikk66(644)

Let L be length and B the breadth.
Perimeter = 2*(L+B) = 84
L+B = 42
If 4 cm is taken from length it becomes L - 4
If 4 cm is added to breadth it becomes B + 4
Since it becomes a square, and all sides of square are equal, we have
L - 4 = B + 4
L = B + 8
Since L + B = 42,
Substituting for L as B + 8
B + 8 + B = 42
2*B = 34
B = 17 cm
Length = B + 8 = 25 cm
Hope you got it :)